OpenStudy (electronia):

3b-9=15 I will medal for a step by step answer! :)

OpenStudy (hhmorris98):

add 9 to each side. Then you have 3b=24. Divide by 3 to get b=8

OpenStudy (cathyangs):

3b - 9 = 15 adding 9 to both sides keeps them equal to each other 3b = 15 + 9 = 24 dividing both sides by 3, b = 8.

OpenStudy (agl202):

Reorder the terms: -9 + 3b = 15 Solving -9 + 3b = 15 Solving for variable 'b'. Move all terms containing b to the left, all other terms to the right. Add '9' to each side of the equation. -9 + 9 + 3b = 15 + 9 Combine like terms: -9 + 9 = 0 0 + 3b = 15 + 9 3b = 15 + 9 Combine like terms: 15 + 9 = 24 3b = 24 Divide each side by '3'. b = 8 Simplifying b = 8

OpenStudy (electronia):

Thank you! :)

OpenStudy (agl202):

Your welcome!